PRECISE DETERMINATION OF DIPOLE MOMENT OF METHANOL BY MICROWAVE STARK SPEC-TROSCOPY

\underline{INDRANATH MUKHOPADHYAY}, Laser Programme, Centre for Advanced Technology, Indore 452 013, India; K.V.L.N. SASTRY, Department of Physics, University of New Brunswick, Fredericton, N.B., E3B 5A3, Canada.

In this work, precise Stark effect measurements have been carried out on several transitions in the first three torsional states of methanol. The Stark shifted transition frequencies for a wide range of steady electric field have been measured with an accuracy of about 10 kHz. Detailed analysis of the data allowed the determination of the dipole moment components in the first three torsional states of the ground vibrational state and the corresponding zero field frequencies have been determined. The present study represents the measurement of the most accurate dipole moment values of methanol.